OpenStudy (anonymous):

Two brothers enter a race with five friends. The racers draw lots to determine their starting positions. What is the probability that the older brother will start in lane 1 with his brother beside him in lane 2?

OpenStudy (kropot72):

Assumin the brothers are the first two to make draws the probability that they draw two specific lanes is given by \[\frac{\left(\begin{matrix}2 \\ 2\end{matrix}\right)\left(\begin{matrix}5 \\ 0\end{matrix}\right)}{\left(\begin{matrix}7 \\ 2\end{matrix}\right)}=\frac{1}{21}\] There are 2! permutations of the brothers' positions. So the probability of a specific permutation is \[\frac{1}{21}\times \frac{1}{2}\]

OpenStudy (kropot72):

Assuming*

OpenStudy (anonymous):

thankss :) why do you multiply the 1/21 by 1/2 ?

OpenStudy (kropot72):

The reason for multiplying 1/21 by 1/2 is as follows: If the brothers are the first two competitors to make draws, the probability that one brother draws lane 1 and the other brother draws lane 2 is 1/21. But the required probability is that the older brother draws lane 1 and the younger brother draws lane 2. There are 2 possible outcomes if one brother draws lane 1 and the other brother draws lane 2: 1. Younger brother draws lane 1 and older brother draws lane 2. 2. Older brother draws lane 1 and younger brother draws lane 2. Only outcome #2 is allowed, therefore 1/21 must be divided by 2 to get the required probability.